OpenStudy (anonymous):

sec[inverse sin (2sqrt (6)/ 11]

zepdrix (zepdrix):

\[\large \sec\left(\color{royalblue}{\arcsin\left(\frac{2\sqrt6}{11}\right)}\right)\]It looks like you're missing a bracket somewhere, does this look correct? The 11 is inside of the arcsine function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

Just draw triangles.

OpenStudy (anonymous):

can i draw triangles for the ones on the unit circle if i cant remember the ordered pairs?

OpenStudy (anonymous):

you can always draw triangles when you have an inverse trig function followed by a trig function.

OpenStudy (anonymous):

ok (:

OpenStudy (anonymous):

i got sqrt 97 for the adj side

OpenStudy (anonymous):

would i get 11/sqrt(97)?